OpenStudy (anonymous):

if x=m^3 and y=2n^2, write the following in terms of m and r. 3x^2y

jimthompson5910 (jim_thompson5910):

I think you meant to say "in terms of m and n" (not m and r)

jimthompson5910 (jim_thompson5910):

am I correct with that assumption?

OpenStudy (anonymous):

the paper says m and r but I thinks its m and n

jimthompson5910 (jim_thompson5910):

yeah I think it's m and n too because r is nowhere to be found

jimthompson5910 (jim_thompson5910):

x=m^3 x^2=(m^3)^2 ... square both sides x^2=m^(3*2) x^2=m^6

jimthompson5910 (jim_thompson5910):

3x^2y 3m^6y ... replace x^2 with m^6 (since x^2 = m^6, found earlier) 3m^6*2n^2 ...replace y with 2n^2 (because it's given that y = 2n^2) 3*2m^6n^2 6m^6n^2 so this means that 3x^2y = 6m^6n^2 when x = m^3 and y = 2n^2

OpenStudy (anonymous):

ohh it makes much more sense, thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome